OpenStudy (anonymous):

Assume that y varies inversely with x. If y = -3 when x = 2.4, what is y when x = 2? -3.6 -0.8 -7.2 4.8

OpenStudy (alexwee123):

y=k/x -3=k/2.4 k=-7.2 y=-7.2/x y=-7.2/2 y=-3.6

OpenStudy (alexwee123):

If y varies inversely with x, then y is proportional to 1/x. This is the same thing as saying the y equal 1/x times some constant. So you have y = k/x

OpenStudy (alexwee123):

y=-3 when x=-2.4 so k=-7.2

OpenStudy (alexwee123):

then find y when x=2

OpenStudy (alexwee123):

x=2.4* sorry i made a typo